# On the Isbell Problem

### v1

@@ Abstract.
Acknowledgments. The first author was supported by the “Programme to support prospective human resources – post Ph.D. candidates” of the Czech Academy of Sciences, project L100192251. The research for this paper was started when the first author was a postdoc at Institute of Mathematics of Czech Academy of Sciences, and concluded after he became a postdoc in the Universidad Autónoma de México (UNAM).
@@
Theorem 1.1 is built via a countable support iteration of proper forcings resulting from a reformulation of Shelah’s forcing from [22]. Theorem 1.2 is a more demanding task and needs the development of new forcing techniques. In all the models we construct, there is no $\mathsf{nwd}$-ultrafilter, and therefore, Theorem 1.2 gives an answer to Problem 3 from [1] of J. Brendle on the consistency of the non-existence of nowhere dense ultrafilters and the continuum being larger than $\omega_{2}$. We produce two models in which all ultrafilters are Tukey top. In both constructions, we first make a forcing extension by an $\omega_{2}$-length countable support iteration of the Sacks forcing. Then we define suborder of carefully chosen countable support product; this is the step in which both constructions differ: one construction is made through a generalization of countable support iterations of strongly proper forcings having one additional property, while the other relies on the existence of square sequences and one theorem of Woodin about the universe $L[S]$, where $S$ is a sequence of ordinals. However, forcings of the kind $\mathbb{P}*\dot{\mathbb{Q}}$, where $\mathbb{P}$ is strongly proper and $\dot{\mathbb{Q}}$ is Cohen preseving, play a central role in both construction. Both constructions produce forcings which preserve $\omega_{1}$, are $\aleph_{2}$-c.c. and Cohen preserving.
@@
The structure of the paper is as follows: section 2 introduces the basic results on the critical ideals. These ideals pose the question on the possibility of non-nowhere dense ultrafilters being Tukey top (or at least non basically generated), which we answer by providing a construction of a basically generated ultrafilter which is not nowhere dense. Section 3 introduces a kind of filters generalizing the notion of saturated filters and which will be of importance for the construction of our forcings from sections 5 and 8. Section 4 is dedicated to the development of the basic forcing $\mathbb{P}_{\kappa}(\mathcal{U})$, its properties and the proof of Theorem 1.1. Section 5 introduces the first version of the poset we use to prove Theorem 1.2 and proves the forcing has the $\aleph_{2}$-c.c. Section 6 introduces an new iteration technique from which the forcing from section 5 is a special case. Section 7 completes the construction of the first model for Theorem 1.2. Section 8 is dedicated to the construction of the second model for Theorem 1.2; this is a more direct solution to the Isbell problem, and the construction is based on the existence of square sequences and one theorem of Woodin, and provides an existence theorem of a reduced product (that is, a suborder of a countable support product) which is interesting by itself. The paper finishes with a section of open questions.
@@
Let $\mathscr{I,J}$ be two ideals. We say that $\mathscr{I}$ is Katětov below $\mathscr{J}$, denote by $\mathscr{I}\leq_{K}\mathscr{J}$, if there is a function $f:\bigcup\mathscr{J}\to\bigcup\mathscr{I}$ such that for all $A\in\mathscr{I}$, $f^{-1}[A]\in\mathscr{J}$.
@@
An ultrafilter $\mathcal{U}$ is basically generated if there is $\mathcal{B}\subseteq\mathcal{U}$ which generates $\mathcal{U}$, and for any $\{X_{n}:n\in\omega\}\subseteq\mathcal{B}$ such that $\lim_{n\in\omega}X_{n}\in\mathcal{B}$, there is $A\in[\omega]^{\omega}$ such that $\bigcup_{n\in\omega}X_{n}\in\mathcal{U}$.
@@
This section introduces some terminology that will be used in sections 5 to 7. Section 4 does not require everything that is presented here, and all that it needs is that whenever $\mathcal{U}$ is an ultrafilter, then Player I has no winning strategy in any of the games defined in 2.3, 2.12, 2.13 and 2.14 below; these are classical and well known results.
@@ II $\tau_{0}(f\upharpoonright 1)$ $\tau_{0}(f\upharpoonright 2)$ $\tau_{0}(f\upharpoonright 3)$ldots
Let $p\in\mathbb{P}_{\kappa}(\mathcal{F})$ be a condition and $\vec{\nu}=\langle\nu_{n}:n\in\omega\rangle\in\mathcal{C}_{\kappa}$. We can assume that $p$ is a simple condition and that for all $n\in\omega$, $dom(\nu_{n})\subseteq dom(p)$. Let $\langle\alpha_{n}:n\in\omega\rangle$ be the canonical enumeration of $dom(p)$. First note that for all $k\in\omega$, there are $m\geq k$ and $l\in\omega$ such that $dom(\nu_{l})\subseteq\{\alpha_{j}:j\in[k,m)\}$. We give a strategy to Player I in the game $\mathcal{G}^{\mathcal{P}}(\mathcal{F})$, and on the side we are going to construct sequences $\langle p_{n}:n\in\omega\rangle,\langle m_{n}:n\in\omega\rangle,\langle k_{n}:n\in\omega\rangle,\langle\vec{E}_{n}^{0},\vec{E}_{n}^{1},E_{n}^{-1}:n\in\omega\rangle,\langle b_{n}:n\in\omega\rangle$
@@
6. A new kind of iterations.
In this section we develop a general framework from which the forcing $\mathbb{Q}_{\kappa}$ is a special case. We define $\varsigma$-iterations and develop the basic properties of this kind of iteration, finishing with the preservation of $\varsigma$-properness along $\varsigma$-iterations.
@@
Let $\mathbb{P}$ be a forcing. We say that $\mathbb{P}$ is strongly proper if for any $\mathcal{M}\prec H(\theta)$ (for big enough $\theta$) such that $\mathbb{P}\in\mathcal{M}$, any $p\in\mathbb{P}\cap\mathcal{M}$ and any family $\vec{\mathcal{D}}=\{D_{n}:n\in\omega\}$ of open dense subsets of $\mathbb{P}\cap\mathcal{M}$, there is a condition $q\leq p$ such that each set $D_{n}$ is predense below $q$. We say that $q$ is a $(\mathbb{P},\mathcal{M},\vec{\mathcal{D}})$-generic condition.
It is well known that countable support iterations of strongly proper forcings are strongly proper (see [21]).
The following lemma is easy to prove, so we omit the proof.
@@
Next we fix some notation to be used along the rest of the section.
@@
    If $f$ is a function and $A$ is a set, then $f*A=\{(a,b)\in f:a\notin A\}$(when taking about iterations * has its usual meaning).
@@
Let $\mathbb{P}$ and $\mathbb{Q}$ be forcings such that $\mathbb{P}\lessdot\mathbb{Q}$ and let $\pi:\mathbb{Q}\to\mathbb{P}$ be a projection. Let $\dot{\mathbb{R}}$ be a $\mathbb{Q}$-name for a forcing and consider $\mathbb{Q}*\dot{\mathbb{R}}$. For $X\subseteq\mathbb{Q}*\dot{\mathbb{R}}$, define the following sets:
  1. (1)
    $X^{\divideontimes}=\{(\dot{r},q):(q,\dot{r})\in X\}$.
  2. (2)
    ${}_{\pi}X=\{(\pi(q),(q,\dot{r})):(q,\dot{r})\in X\}$.
  3. (3)
    ${{}_{\pi}X}^{\divideontimes}=(_{\pi}X)^{\divideontimes}=\{((q,\dot{r}),\pi(q)):(q,\dot{r})\in X\}$.
Also, whenever $\pi$ is a proyection and $\dot{x}$ a $\mathbb{Q}$-name, we denote by $\pi(\dot{x})$ the corresponding $\mathbb{P}$-name for the $[\mathbb{Q}:\mathbb{P}]$-name for $\dot{x}$.
The next lemma are easy remarks.
Lemma 6.4.
Let $\mathbb{P}$ and $\mathbb{Q}$ be forcings such that $\mathbb{P}\lessdot\mathbb{Q}$. For any $X\subseteq\mathbb{Q}*\dot{\mathbb{R}}$:
  1. (1)
    $X^{\divideontimes}$ is a $\mathbb{Q}$-name for a subset of $\dot{\mathbb{R}}$.
  2. (2)
    ${}_{\pi}X^{\divideontimes}$ is a $\mathbb{P}$-name for a subset of $\mathbb{Q}*\dot{\mathbb{R}}$.
  3. (3)
    If $X$ is open dense, then ${}_{\pi}X^{\divideontimes}$ is a $\mathbb{P}$-name for an open dense subset of $[\mathbb{Q}*\dot{\mathbb{R}}:\mathbb{P}]$.
Definition 6.5.
Let $\mathbb{P,Q}$ be forcings such that $\mathbb{P}\lessdot\mathbb{Q}$ and $\mathbb{Q}\to\mathbb{P}$ the corresponding projection, and $\dot{\mathbb{R}}$ a $\mathbb{P}$-name for a forcing. Define the $\mathbb{P}$-restricted iteration of $\mathbb{Q}$ and $\dot{\mathbb{R}}$ as: $(p,\dot{r})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}$ if and only if:
  1. (1)
    $(q,\dot{r})\in\mathbb{Q}\times dom(\dot{\mathbb{R}})$.
  2. (2)
    $\pi(q)\Vdash\dot{r}\in\dot{\mathbb{R}}$.
@@
Let $\mathbb{P}$ and $\mathbb{Q}$ be complete Boolean algebras such that $\mathbb{P}\lessdot\mathbb{Q}$, and let $\pi$ be the projection from $\mathbb{Q}$ to $\mathbb{P}$. We say that $(\mathbb{P},\mathbb{Q})$ is $\varsigma$-proper if for any countable $\mathcal{M}\prec H(\theta)$ such that $\mathbb{P},\mathbb{Q}\in\mathcal{M}$, any $\mathbb{P}$-name $\dot{\mathbb{R}}\in\mathcal{M}$ for a forcing, any $q\in\mathcal{M}\cap\mathbb{Q}$ and $\langle D_{n}:n\in\omega\rangle$ such that for each $n\in\omega$:
@@
    $\pi(q_{0})\Vdash\dot{E}_{n}\text{ is an open dense subset of $\dot{\mathbb{R}}\cap\mathcal{M}[G]$}$.
@@
where $H$ is $\mathbb{Q}$-generic and $G$ is the $\pi$-projection of $H$.
We say that $q_{0}$ is a weakly $(\mathbb{P},\mathbb{Q},\mathcal{M})$-generic condition for $\{D_{n}:n\in\omega\}$ and $\dot{\mathbb{R}}$.
While this definition may seem quite restrictive, the next lemmas give a simpler description of the notion which allow us to say that such pairs of forcings are not scarce.
Proposition 6.7.
If $(\mathbb{P},\mathbb{Q})$ is $\varsigma$-proper, then $\mathbb{Q}$ is strongly proper.
@@
Let $\mathcal{M}\prec H(\theta)$ be countable such that $\mathbb{P},\mathbb{Q}\in\mathcal{M}$ and fix $q\in\mathbb{Q}\cap\mathcal{M}$. Let us work with $\mathbb{Q}*\check{\mathbb{C}}$, where $\check{\mathbb{C}}$ is the check name for the Cohen forcing, so we can assume $\mathbb{Q}*\check{\mathbb{C}}$ is $\mathbb{Q}\times\mathbb{C}$. Let $\{D_{n}:n\in\omega\}$ be a family of open dense subsets of $\mathbb{Q}\cap\mathcal{M}$. Let $F_{0}\subseteq\mathbb{Q}\times\mathbb{C}\cap\mathcal{M}$ be an open dense subset of $\mathcal{M}\cap\mathbb{Q}\times\mathbb{C}$. Define, for each $n\in\omega$ a new open dense subset of $\mathcal{M}\cap\mathbb{Q}\times{\mathbb{C}}$:
$\overline{D}_{n}=\{(a,{b})\in F_{0}:a\in D_{n}\}$
$\overline{D}_{n}$ is open dense in $\mathbb{Q}\times{\mathbb{C}}\cap\mathcal{M}$: pick $(p,{r})\in\mathbb{Q}\times{\mathbb{C}}\cap\mathcal{M}$, since $F_{0}$ is dense, there is $(a,{b})\in F_{0}$ such that $(a,{b})\leq(p,{r})$, since $D_{n}$ is dense, we find $c\in D_{n}$ such that $c\leq a$, so $(c,{b})\leq(a,{b})$, and since $F_{0}$ is open, $(c,{b})\in F_{0}$, thus, $(c,{b})\in\overline{D}_{n}$; this proves $\overline{D}_{n}$ is dense. Now let $(p,{r})\in\overline{D}_{n}$ and $(a,{b})\leq(p,{r})$ such that $(a,{b})\in\mathbb{Q}\times{\mathbb{C}}\cap\mathcal{M}$. By definition of $\overline{D}_{n}$, $(p,{r})\in F_{0}$ and $p\in D_{n}$, so $a\in D_{n}$ and $(a,{b})\in F_{0}$, which implies $(a,{b})\in\overline{D}_{n}$; thus, $\overline{D}_{n}$ is open.
Now extend $q$ to a condition $(\mathbb{P},\mathbb{Q},\mathcal{M})$-generic $q_{0}$ for $\{\overline{D}_{n}:n\in\omega\}$ and $\mathbb{S}$. Then $q_{0}$ forces $\overline{D}_{n}^{\divideontimes}$ is an open dense subset of ${\mathbb{C}}\cap\mathcal{M}[\dot{H}]$, where $\dot{H}$ is a $\mathbb{Q}$-generic filter such that $q_{0}\in H$. Note that this implies that each $D_{n}$ is predense below $q_{0}$: otherwise, there are $q^{\prime}\leq q_{0}$ and $m_{0}\in\omega$ such that $q^{\prime}$ incompatible with each element of $D_{m_{0}}$, which implies that for each $(s,\dot{r})\in\overline{D}_{n}\cap\mathcal{M}$, $s$ is incompatible with $q^{\prime}$, which implies that $q^{\prime}$ forces $\overline{D}_{n}^{\divideontimes}=\emptyset$, a contradiction. ∎
Lemma 6.8.
If $(\mathbb{P},\mathbb{Q})$ is $\varsigma$-proper, then $\mathbb{P}$ forces $[\mathbb{Q}:\mathbb{P}]$ is Cohen preserving. A fortiori, $[\mathbb{Q}:\mathbb{P}]$ is $\omega^{\omega}$-bounding over $V^{\mathbb{P}}$.
Proof.
Let $\dot{D}$ be a $\mathbb{P}$-name for a $[\mathbb{Q}:\mathbb{P}]$-name and $(p_{0},\dot{q}_{0})\in\mathbb{P}*[\mathbb{Q}:\mathbb{P}]$ a condition such that $p_{0}$ forces $\dot{q}_{0}$ forces $\dot{D}$ to be an open dense subset of $\mathbb{C}$. Then we can assume $\dot{D}$ is a $\mathbb{P}*[\mathbb{Q}:\mathbb{P}]$-name for an open dense subset of $\mathbb{C}$. Let $\mathcal{M}\prec H(\theta)$ be countable and such that $(p_{0},\dot{q}_{0}),\mathbb{P},\dot{D}\in\mathcal{M}$. Now define a set $\overline{D}$ as $((a,\dot{b}),s)\in\overline{D}$ if and only if:
  1. (1)
    $((a,\dot{b}),{s})\in(\mathbb{P}*[\mathbb{Q}:\mathbb{P}])\times{\mathbb{C}}\cap\mathcal{M}$.
  2. (2)
    There are $(\dot{r},(c,\dot{d}))\in\dot{D}\cap\mathcal{M}$ and $t\in\mathbb{C}$ such that:
    1. (a)
      $t\subseteq s$.
    2. (b)
      $(a,\dot{b})\leq(c,\dot{d})$.
    3. (c)
      $(a,\dot{b})\Vdash\dot{r}=t$.
Let us see that $\overline{D}$ is an open dense subset of $\mathbb{P}*[\mathbb{Q}:\mathbb{P}]\times\mathbb{C}\cap\mathcal{M}$. Let $((a,\dot{b}),{s})\in(\mathbb{P}*[\mathbb{Q}:\mathbb{P}]\times{\mathbb{C}})\cap\mathcal{M}$ be arbitrary. Since $\dot{D}$ is forced to be dense, we find, inside $\mathcal{M}$, $(c_{1},\dot{d}_{1})\leq(a,\dot{b})$ and $(\dot{r},(c_{0},\dot{d}_{0}))\in\dot{D}$ such that $(c_{1},\dot{d}_{1})\leq(c_{0},\dot{d}_{0})$ and $(c_{1},\dot{d}_{1})\Vdash\check{s}\subseteq\dot{r}$, and we can extend $(c_{1},\dot{d}_{1})$ to a condition $(c_{2},\dot{d}_{2})\in\mathbb{P}*[\mathbb{Q}:\mathbb{P}]\cap\mathcal{M}$ and find $r_{0}\in\mathbb{C}$ such that $(c_{2},\dot{d}_{2})\Vdash\dot{r}=r_{0}$. Then we have $s\subseteq r_{0}$. Now, we have $(\dot{r},(c_{0},\dot{d}_{0}))\in\dot{D}\cap\mathcal{M}$, $r_{0}\in\mathbb{C}$, $(c_{2},\dot{d}_{2})\leq(c_{0},\dot{d}_{0}),(a,\dot{b})$, and $(c_{2},\dot{d}_{2})\Vdash\dot{r}=r_{0}$, so $((c_{2},\dot{d}_{2}),{r}_{0})\in\overline{D}$ and extends $((a,\dot{b}),{s})$. That $\overline{D}$ is open also follows easily from the definition.
Let us now extend $(p_{0},\dot{q}_{0})$ to a $(\mathbb{P},\mathbb{P}*[\mathbb{Q}:\mathbb{P}],\mathcal{M})$-generic condition $(p,\dot{q})$ for $\overline{D}$ and $\mathbb{C}$, and let $\dot{E}$ be the corresponding $\mathbb{P}$-name for an dense open subset of $\mathbb{C}\cap\mathcal{M}[G]$ which is forced to be a subset of $(\overline{D})^{\divideontimes}$. Let $G*H$ be a $\mathbb{P}*[\mathbb{Q}:\mathbb{P}]$-generic filter such that $(p,\dot{q})\in\ G*H$. Then, in $V[G]$, we have $\dot{E}\subseteq\mathbb{C}$ is open dense, and in $V[G*H]$ $\dot{E}\subseteq\overline{D}^{\divideontimes}$, so, for each $s\in\dot{E}$, there is $(a,\dot{b})\in G*H$ and ${r}$ such that $((a,b),{r})\in\overline{D}$ and $s=r$. By definition of $\overline{D}$, we find $(\dot{t},(c,\dot{d}))\in\dot{D}$ and $\sigma\in\mathbb{C}$ such that:
$(a,\dot{b})\leq(c,\dot{d})$
$(a,\dot{b})\Vdash\dot{t}=\sigma\subseteq r$
therefore, $(c,\dot{d})\in G*H$, which implies $\sigma\in\dot{D}_{G*H}$, from which follows $r\in\dot{D}_{G*H}$, that is $s\in\dot{D}_{G*H}$. Thus, we have $\dot{E}\subseteq\dot{D}_{G*H}$. It follows that $\mathbb{P}$ forces $[\mathbb{Q}:\mathbb{P}]$ is Cohen preserving. ∎
Lemma 6.9.
Let $\mathbb{Q}$ be a strongly proper forcing, $\mathbb{P}\lessdot\mathbb{Q}$, and $\mathcal{M}\prec H(\theta)$ countable such that $\mathbb{P},\mathbb{Q}\in\mathcal{M}$. Let $\dot{\mathbb{R}}\in\mathcal{M}$ be a $\mathbb{P}$-name for a forcing, so $\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\in\mathcal{M}$, and $\{D_{n}:n\in\omega\}$ a family of open dense subsets of $\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M}$. Fix $q\in\mathbb{Q}\cap\mathcal{M}$. Then there is $q_{0}\leq q$ which forces that for each $n\in\omega$, $D_{n}^{\divideontimes}$ is an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\dot{G}]$.
Proof.
@@
We need to define several open dense subsets of $\mathbb{Q}\cap\mathcal{M}$:

### v2

@@ Abstract.
2020 Mathematics Subject Classification:
03E35, 03E04, 03E05
@@
Theorem 1.1 is built via a countable support iteration of proper forcings resulting from a reformulation of Shelah’s forcing from [22]. Theorem 1.2 is a more demanding task and needs the development of new forcing techniques. In all the models we construct, there is no $\mathsf{nwd}$-ultrafilter, and therefore, Theorem 1.2 gives an answer to Problem 3 from [1] of J. Brendle on the consistency of the non-existence of nowhere dense ultrafilters and the continuum being larger than $\omega_{2}$. We produce two models in which all ultrafilters are Tukey top. In both constructions, we first make a forcing extension by an $\omega_{2}$-length countable support iteration of the Sacks forcing. Then we define suborder of carefully chosen countable support product; this is the step in which both constructions differ: one construction is made through a generalization of countable support iterations of strongly proper forcings [^1] Here we refer to Shelah’s strongly proper forcing notion, instead of the more well known Mitchell’s strongly proper forcing notion; see Definition 6.1 and the paragraph before it. having one additional property, while the other relies on the existence of square sequences and one theorem of Woodin about the universe $L[S]$, where $S$ is a sequence of ordinals. However, forcings of the kind $\mathbb{P}*\dot{\mathbb{Q}}$, where $\mathbb{P}$ is strongly proper and $\dot{\mathbb{Q}}$ is Cohen preseving, play a central role in both construction. Both constructions produce forcings which preserve $\omega_{1}$, are $\aleph_{2}$-c.c. and Cohen preserving.
@@
The structure of the paper is as follows: section 2 introduces the basic results on the critical ideals. These ideals pose the question on the possibility of non-nowhere dense ultrafilters being Tukey top (or at least non basically generated), which we answer by providing a construction of a basically generated ultrafilter which is not nowhere dense. Section 3 introduces a kind of filters generalizing the notion of saturated filters and which will be of importance for the construction of our forcings from sections 4, 5 and 10. Section 4 is dedicated to the development of the basic forcing $\mathbb{P}_{\kappa}(\mathcal{U})$, its properties and the proof of Theorem 1.1. Section 5 introduces the first version of the poset we use to prove Theorem 1.2 and proves the forcing has the $\aleph_{2}$-c.c. Section 6 introduces the notion of Cohen properness which is the central property of our iterations, and prove some equivalent formulations of it. In section 7 we introduce one way to iteration Cohen proper forcings and in section 8 we prove these iterations preserve the notion of Cohen properness. Section 9 completes the construction of the first model for Theorem 1.2. Section 10 is dedicated to the construction of the second model for Theorem 1.2; this construction is based on the existence of square sequences and one theorem of Woodin, and provides an existence theorem of a reduced product (that is, a suborder of a countable support product) which is interesting by itself. The paper finishes with a section of remarks and open questions.
@@
Let $\mathscr{I,J}$ be two ideals. We say that $\mathscr{I}$ is Katětov below $\mathscr{J}$, denoted by $\mathscr{I}\leq_{K}\mathscr{J}$, if there is a function $f:\bigcup\mathscr{J}\to\bigcup\mathscr{I}$ such that for all $A\in\mathscr{I}$, $f^{-1}[A]\in\mathscr{J}$.
@@
An ultrafilter $\mathcal{U}$ is basically generated if there is $\mathcal{B}\subseteq\mathcal{U}$ which generates $\mathcal{U}$, and for any $\{X_{n}:n\in\omega\}\subseteq\mathcal{B}$ such that $\lim_{n\in\omega}X_{n}\in\mathcal{B}$, there is $A\in[\omega]^{\omega}$ such that $\bigcap_{n\in\omega}X_{n}\in\mathcal{U}$.
@@
This section introduces some terminology that will be used in sections 5 to 9. Section 4 does not require everything that is presented here, and all that it needs is that whenever $\mathcal{U}$ is an ultrafilter, then Player I has no winning strategy in any of the games defined in 2.3, 2.12, 2.13 and 2.14 below; these are classical and well known results.
@@ II $\tau_{0}(f\upharpoonright 1)$ $\tau_{0}(f\upharpoonright 2)$ $\tau_{0}(f\upharpoonright 3)$ldots
Let $p\in\mathbb{P}_{\kappa}(\mathcal{F})$ be a condition and $\vec{\nu}=\langle\nu_{n}:n\in\omega\rangle\in\mathcal{C}_{\kappa}$. We can assume that $p$ is a simple condition and that for all $n\in\omega$, $dom(\nu_{n})\subseteq dom(p)$. Let $\langle\alpha_{n}:n\in\omega\rangle$ be the canonical enumeration of $dom(p)$. First note that for all $k\in\omega$, there are $m\geq k$ and $l\in\omega$ such that $dom(\nu_{l})\subseteq\{\alpha_{j}:j\in[k,m)\}$. We give a strategy to Player I in the game $\mathcal{G}^{\mathcal{P}}(\mathcal{F})$, and on the side we are going to construct sequences $\langle p_{n}:n\in\omega\rangle,\langle m_{n}:n\in\omega\rangle,\langle k_{n}:n\in\omega\rangle,\langle\vec{E}_{n}^{0},\vec{E}_{n}^{1},E_{n}^{-1}:n\in\omega\rangle,\langle b_{n}:n\in\omega\rangle$.
@@
6. Cohen proper forcing.
In this section and the next one we develop a general framework from which the forcing $\mathbb{Q}_{\kappa}$ is a special case. Here we define the notion of Cohen properness and derive some equivalent formulations enclosed in Theorem 6.8. The reader that wish to jump to the iteration and preservation theorem can take the equivalence given in Theorem 6.8 clause (4) as the definition of Cohen proper forcing and go directly to the next section (in this case the reader will also need definitions 6.3 and 6.4 below).
The notion stated in the next definition was named strongly proper by Shelah. However, there is also a strongly proper forcing notion due to Mitchell. It seems that Mitchell’s notion of strongly proper forcing is more well known, so, to avoid any possible confusion and fully distinguish the two notions, we have decided to rename Shelah’s notion as “$\sigma$-proper”. More explicitly:
@@
Let $\mathbb{P}$ be a forcing. We say that $\mathbb{P}$ is $\sigma$-proper if for any $\mathcal{M}\prec H(\theta)$ (for big enough $\theta$) such that $\mathbb{P}\in\mathcal{M}$, any $p\in\mathbb{P}\cap\mathcal{M}$ and any family $\vec{\mathcal{D}}=\{D_{n}:n\in\omega\}$ of open dense subsets of $\mathbb{P}\cap\mathcal{M}$, there is a condition $q\leq p$ such that each set $D_{n}$ is predense below $q$. We say that $q$ is a $(\mathbb{P},\mathcal{M},\vec{\mathcal{D}})$-generic condition.
It is well known that countable support iterations of $\sigma$-proper forcings are $\sigma$-proper (strongly proper, under Shelah’s naming) (see [21]).
The following lemma is easy to prove, so we omit the argument.
@@
Next we fix some notation to be used along the rest of the paper.
@@
    If $f$ is a function and $A$ is a set, then $f*A=\{(a,b)\in f:a\notin A\}$(when taking about iterations * keeps its usual meaning, the context should be enough to differentiate between the two meanings).
@@
Let $\mathbb{P,Q}$ be forcings such that $\mathbb{P}\lessdot\mathbb{Q}$, and let $\pi:\mathbb{Q}\to\mathbb{P}$ be a regular projection projection, $\dot{\mathbb{R}}$ a $\mathbb{P}$-name for a forcing and $\mathbb{P}*\dot{\mathbb{R}}$ the two step iteration of $\mathbb{P}$ and $\dot{\mathbb{R}}$. Define the $\mathbb{P}$-restricted iteration of $\mathbb{Q}$ and $\dot{\mathbb{R}}$ as: $(q,\dot{r})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}$ if and only if:
$(\pi(q),\dot{r})\in\mathbb{P}*\dot{\mathbb{R}}.$
@@
Let $\mathbb{P}$ and $\mathbb{Q}$ be forcings such that $\mathbb{P}\lessdot\mathbb{Q}$ and let $\pi:\mathbb{Q}\to\mathbb{P}$ be a projection. Let $\dot{\mathbb{R}}$ be a $\mathbb{Q}$-name for a forcing and consider $\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}$. For $X\subseteq\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}$, define the set
$X^{\divideontimes}=\{(\dot{r},q):(q,\dot{r})\in X\}$
Note that $X^{\divideontimes}$ is a $\mathbb{Q}$-name for a subset of $\dot{\mathbb{R}}$.
Definition 6.5.
Let $\mathbb{P}$ and $\mathbb{Q}$ be complete Boolean algebras such that $\mathbb{P}\lessdot\mathbb{Q}$, and let $\pi$ be the projection from $\mathbb{Q}$ to $\mathbb{P}$. We say that $(\mathbb{P},\mathbb{Q})$ is Cohen proper if for any countable $\mathcal{M}\prec H(\theta)$ such that $\mathbb{P},\mathbb{Q}\in\mathcal{M}$, any $\mathbb{P}$-name $\dot{\mathbb{R}}\in\mathcal{M}$ for a forcing, any $q\in\mathcal{M}\cap\mathbb{Q}$ and $\langle D_{n}:n\in\omega\rangle$ such that for each $n\in\omega$:
@@
    $\pi(q_{0})\Vdash\dot{E}_{n}\text{ is an open dense subset of $\dot{\mathbb{R}}\cap\mathcal{M}[\pi[\dot{G}]]$}$.
  3. (3)
    $q_{0}\Vdash D_{n}^{\divideontimes}\subseteq\dot{\mathbb{R}}\cap\mathcal{M}[\dot{G}]\text{ is open dense}$.
@@
where $G$ is $\mathbb{Q}$-generic.
We say that $q_{0}$ is a $(\mathbb{P},\mathbb{Q},\mathcal{M})$-generic condition for $\{D_{n}:n\in\omega\}$ and $\dot{\mathbb{R}}$.
Lemma 6.6.
Let $\mathbb{Q}$ be a $\sigma$-proper forcing, $\mathbb{P}\lessdot\mathbb{Q}$ a regular suborder, and $\mathcal{M}\prec H(\theta)$ countable such that $\mathbb{P},\mathbb{Q}\in\mathcal{M}$. Let $\dot{\mathbb{R}}\in\mathcal{M}$ be a $\mathbb{P}$-name for a forcing, so $\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\in\mathcal{M}$, and $\{D_{n}:n\in\omega\}$ a family of open dense subsets of $\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M}$. Fix $q\in\mathbb{Q}\cap\mathcal{M}$. Then there is $q_{0}\leq q$ which forces that for each $n\in\omega$, $D_{n}^{\divideontimes}$ is an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\dot{G}]$.
@@
To simplify notation, given a condition $a\in\mathbb{Q}$, we write $a^{\perp}$ to denote the set of conditions in $\mathbb{Q}$ which are incompatible with $a$. We need to define several open dense subsets of $\mathbb{Q}\cap\mathcal{M}$: